# Synergies Between Federated Learning and O-RAN: Towards an Elastic Architecture for Multiple Distributed Machine Learning Services

### v4

@@ Abstract
Federated learning (FL) is a popular distributed machine learning (ML) technique. However, implementation of FL over 5G-and-beyond wireless networks faces key challenges caused by (i) dynamics of the wireless network conditions and (ii) the coexistence of multiple FL-services in the system, which are not jointly considered in prior works. We first take a closer look into these challenges and unveil nuanced phenomena called over-/under-provisioning of resources and perspective-driven load balancing. We then take the first steps towards addressing these phenomena by proposing a novel distributed ML architecture called elastic FL (EFL). EFL unleashes the full potential of Open RAN (O-RAN) systems and introduces an elastic resource provisioning methodology to execute FL-services. It further constitutes a multi-time-scale FL management system that introduces three dedicated network control functionalities tailored for FL-services, including (i) non-real-time (non-RT) system descriptor, which trains ML-based applications to predicted both system and FL-related dynamics and parameters; (ii) near-RT FL controller, which handles O-RAN slicing and mobility management for the seamless execution of FL-services; (iii) FL MAC scheduler, which conducts real-time resource allocation to the end clients of various FL-services. We finally prototype EFL to demonstrate its potential in improving the performance of FL-services.
@@ I Introduction
Federated learning (FL) has attracted tremendous attention for executing data-intensive Internet-of-Things applications (e.g., autonomous driving), where data is distributedly collected at edge devices [1]. This distributed machine learning (ML) approach is an alternative to centralized ML methods that require transferring distributed data to cloud servers, which often cause a high communication overhead and concerns regarding the exposure of raw data to untrusted parties. FL performs ML training through the repetition of multiple global training rounds, with each round consisting of two key steps: (i) using their local data, clients/devices perform local model training, e.g., via several local iterations of stochastic gradient decent (SGD), and transfer their local models to a server, and (ii) the server aggregates all the received models (e.g., via averaging) to a global model and then broadcasts it back to the clients to initiate the next global training round.
@@ I-A Motivation
Despite its tremendous potential, FL is anticipated to be deployed on resource-constrained wireless devices such as smartphones and IoT sensors [1], which is challenging due to the increasing size of modern ML models that demand substantial computational power for local model training and incur high communication overheads during the transfer from devices to the central server. To this end, research – see [1] and the references therein – has focused on (i) tuning the number of local SGD iterations and frequency/period of client-to-server communications; (ii) model compression, gradient sparsification, or selective parameter sharing; (iii) efficient client selection/recruitment. Despite the significance of these methods, there are two key underlying assumptions made in all of them, limiting their applicability in real-world scenarios:
(A1) Snapshot-Based Analysis: These methods often assume that network and client conditions, such as client location and client-to-server channels, remain fixed during each global training round of FL, and thus they conduct snapshot-based analysis of the network. Afterwards, in each snapshot – often presumed to be the condition of the network at the beginning of each global training round – they make control decisions (e.g., spectrum allocation and client recruitment). However, in real-world scenarios, such analysis may quickly become sub-optimal, as the network condition may vary over time due to factors like client mobility and evolving user datasets.
(A2) Isolated Execution of a Single FL-Service: The existing methods primarily focus on executing a single FL-service. However, in large-scale networks, multiple FL-services may simultaneously compete for wireless resources and client recruitment (e.g., Google may run an FL-service for next-word prediction, while Apple does so for face recognition). Despite its importance, only a few studies have explored this topic [2, 3], all of which consider snapshot-based analysis suffering from the above limitations. The coexistence of multiple FL-services further introduces a new dimension of heterogeneity, which we call heterogeneity of interest. This heterogeneity reflects the diverse interests of FL-services in recruiting clients with different (i) datasets, (ii) data quality, (iii) channel conditions, and (iv) processing capabilities. These preferences are shaped by quality-of-service (QoS) requirements of each FL-service (e.g., target model accuracy and latency sensitivity) and its budget to recruit clients and acquire wireless resources.
@@
(2) Perspective-Driven Load Balancing: Load balancing refers to distributing the network traffic, preventing congestion across the network. The coexistence of multiple FL-services (relaxation of (A2)) calls for a unique load balancing strategy that considers a nuanced phenomenon which we call perspective-driven client distribution. This phenomenon captures the fact that the distribution of FL clients across different network regions varies depending on the perspective from which clients are evaluated/perceived. For example, regions with advanced connectivity infrastructure, such as urban areas, may have a higher density of clients suitable for latency-sensitive FL-services. Given the perspective-driven client distribution, in FL, load balancing must account for the overlap of the FL-services’ objectives. For instance, consider two FL-services that need clients from similar regions of the network (e.g., pedestrian detection and line tracking FL-services both executed over smart cars in an urban area). This situation leads to an accumulation of FL-services in that region, causing network traffic congestion.
@@ I-B Migration from Per-User to Service-Level Objectives
Simultaneous addressing of the above two challenges calls for client mobility management (i.e., handover design for load balancing) and dynamic resource allocation solutions, which are the key network control functionalities of radio access networks (RANs) [4, 5]. However, the existing network control functionalities are often developed for native wireless services (e.g., high data-rate Internet provisioning) and primarily focus on per-user QoS requirements, including energy consumption and communication latency. Instead, in FL, addressing the aforementioned two challenges should consider collective service-level objectives/characteristics, such as the prediction accuracy of the FL models and client recruitment costs, alongside QoS metrics like overall energy consumption and model training latency across FL-services. Additionally, effective load balancing in FL must ensure an optimal allocation of clients to FL-services and wireless resources to the clients, while considering the innate differences among FL-services. For example, consider an FL-service for monitoring industrial equipments utilized during business hours coexisting with an FL-service focused on air quality monitoring. Equipment monitoring might require rapid ML model training to predict failures, while environmental sensing may tolerate longer ML training delays. In this case, the dynamic network control functionalities should prioritize allocating network resources to industrial equipment monitoring during business hours and then shift resources towards environmental sensing.
@@
Figure 1: O-RAN architecture (all abbreviations follow standard 3GPP terminologies [6], while FLS stands for FL-service). O-RAN components interact via standard open interfaces (E2, F1, open fronthaul, A1, and O1 in Fig. 1), facilitating interoperability between network elements from different manufacturers.
@@
These considerations necessitate a revamp of the existing network control functionalities to facilitate the coexistence of multiple FL-services within a dynamic network – an approach absent in the literature. This naturally calls for a migration from traditional RANs (e.g., distributed RANs in 4G), operating with black-box network control functionalities applied to all network services [4]. To overcome this limitation and develop dedicated network control functionalities for FL-services, we leverage virtualization, intelligence, and programmability features of an emerging RAN technology used in 5G-and-beyond networks, called Open RAN.
I-C Open RAN (O-RAN) Architecture
5G-and-beyond networks are expected to support diverse services, including ultra-reliable low latency communications and massive machine-type communications [4]. They also are envisioned to provide services for various verticals – groups of companies requiring similar services (e.g., industrial factories) – operated by distinct virtual network operators. This versatile service provisioning is enabled by the use of intelligent and programmable O-RANs. O-RAN (Fig. 1) transitions from rigid cellular to multi-vendor and data-driven networks by integrating the concepts of intelligence, virtualization (RAN slicing), and programmable white-box hardware (as opposed to the black-box hardware) [6]. In O-RAN, 3GPP stack functionalities are disaggregated into (i) radio unit (O-RU), (ii) distributed unit (O-DU), and (iii) centralized unit (O-CU) [6].
@@ I-C Open RAN (O-RAN) Architecture
O-RAN introduces RAN intelligent controllers (RICs), i.e., non-real-time (non-RT) RIC and near-RT RICs, which utilize AI/ML algorithms to manage RAN operations at different time-scales [6, 7]. It further supports multiple services with diverse QoS requirements via RAN slicing, which partitions RAN resources into isolated RAN slices, each leased to a virtual network operator [4]. Unlike traditional RANs, O-RAN fosters dedicated functionalities for each type of service, a feature referred to as programmability. For example, in contrast to the black-box MAC scheduler used in 4G that serves all services uniformly, O-RAN enables the implementation of a dedicated MAC scheduler for each service.
@@
In this work, we aim to introduce FL-services as a new vertical in 5G-and-beyond networks. Specifically, we exploit O-RAN’s unique features, i.e., virtualization, RAN slicing, and programmability and develop a novel intelligent FL architecture, called elastic FL (EFL). In EFL, we introduce dedicated network control functionalities to orchestrate multiple FL-services over a shared dynamic wireless network.
I-D Contributions
@@
  • In EFL, we leverage the intelligence and programmability features of O-RAN to implement dedicated dynamic network control functionalities for FL-services across three time scales – non-RT, near-RT, and RT (relaxation of (A1)).
@@ II EFL: Elastic FL
In the following, we introduce the EFL framework, which will be implemented over O-RAN.
II-A Integration of EFL in O-RAN
Figure 2: The characteristics and nuances of the three main components of EFL: (i) non-RT system descriptor, (ii) near-RT FL controller, (iii) FL MAC scheduler, as well as their interactions.
To integrate EFL in O-RAN, we first focus on intelligent network control in O-RAN. O-RAN supports intelligent control across three levels of agility: (i) Non-RT RIC operates on a timescale of seconds to minutes, establishing network management policies and hosting applications, known as rApps, which handle long-term functionalities like training AI models for network management. These trained AI models will be packaged into applications, referred to by xApps, and deployed on near-RT RIC. (ii) Near-RT RIC, operating on timescales of 10 milliseconds to seconds, leverages xApps to forecast traffic flow and network dynamics. Using these forecasts, near-RT RIC manages load balancing, network slicing, and mobility management. (iii) Finally, the MAC Scheduler, located at the O-DUs, operates in sub-10-millisecond intervals, focusing on real-time resource allocation to the end users.
EFL exploits O-RAN programmability and envisions dedicated functionalities at each of the aforementioned control layers to manage the concurrent execution of multiple FL-services in a dynamic wireless environment (see Fig. 2). We refer to these functionalities as: (i) Non-RT System Descriptor, located at the non-RT RIC; (ii) Near-RT FL Controller, deployed at the near-RT RIC; (iii) FL MAC Scheduler, deployed at the O-DUs. We next design these functionalities to consider the collective service-level performance requirements, such as FL model accuracy across all FL-services.
II-B Non-RT System Descriptor
@@ II-C Near-RT FL Controller
II-C1 FL slice creation
EFL enables the coexistence of multiple FL-services by creating a virtual RAN slice for each FL-service. In particular, using eApps, the near-RT FL controller first predicts the FL system states (e.g., client congestion at O-RUs). These predictions are derived from both current and historical data provided by the non-RT system descriptor. Based on these predictions, the near-RT FL controller then sends a slice creation request for each FL-service to the near-RT RIC, which is responsible for RAN slicing. This request specifies (i) the required resources, such as bandwidth/spectrum, and (ii) the QoS requirements for each FL-service, including delay constraints and priority for accessing O-RAN resources. The slice is then provisioned with the necessary wireless resources (e.g., spectrum) and sent to O-DUs where FL MAC schedulers perform real-time resource allocation for the clients of each FL-service.
@@ II-C1 FL slice creation
Figure 3: Mobility pattern prediction of a client through an eApp (left plot) and the training loss of the eApp (right plot). The eApp can estimate the location of the client, which is then used to measure the client channel gains and their data rates to the O-RUs. These estimations are used in near-RT FL controller to obtain the bandwidth allocation to different RAN slices. Further, they are used in the FL MAC scheduler to conduct resource allocation to the clients.
II-C2 Elastic slice scaling down/up
@@
II-C3 Client mobility management
@@
Figure 4: The number of successful clients and its impact on the model prediction of FL-services under both baseline methods and our proposed EFL method. Our proposed method (depicted on the right side of both plots) outperforms the baseline methods in terms of the number of successful clients and the model prediction performance of FL-services.
II-D FL MAC Scheduler
EFL introduces a new dimension to MAC schedulers, enabling client-level, real-time resource allocation for FL clients. This addresses resource over/under-provisioning at a shorter time-scale compared to near-RT FL controller, tailored to service-level performance metrics of FL-services, such as delay constraints for receiving local models from clients. Specifically, for each RAN slice, we equip EFL with a dedicated FL MAC scheduler located at O-DUs to allocate wireless resources to its clients.

### v6

# Elastic Federated Learning over Open Radio Access Network (O-RAN) for Concurrent Execution of Multiple Distributed Learning Tasks

@@ Abstract
Federated learning (FL) is a popular distributed machine learning (ML) technique in Internet of Things (IoT) networks, where resource-constrained devices collaboratively train ML models while preserving data privacy. However, implementation of FL over 5G-and-beyond wireless networks faces key challenges caused by (i) dynamics of the wireless network conditions and (ii) the coexistence of multiple FL-services in the system. In this paper, we unveil two key phenomena that arise from these challenges: over/under-provisioning of resources and perspective-driven load balancing, both of which significantly impact FL performance in IoT environments. We take the first steps towards addressing these phenomena by proposing a novel distributed ML architecture called elastic FL (EFL). EFL unleashes the full potential of Open RAN (O-RAN) systems and introduces an elastic resource provisioning methodology to execute FL-services. It further constitutes a multi-time-scale FL management system that introduces three dedicated network control functionalities tailored for FL-services, including (i) non-real-time (non-RT) system descriptor, which trains ML-based applications to predict both system and FL-related dynamics and parameters; (ii) near-RT FL controller, which handles O-RAN slicing and mobility management for the seamless execution of FL-services; (iii) FL MAC scheduler, which conducts real-time resource allocation to the end clients of various FL-services. We finally prototype EFL to demonstrate its potential in improving the performance of FL-services. Our implementations are publicly available at the following GitHub repository: https://github.com/payamsiabd/Elastic_FL_Over_O_RAN.git
@@ I Introduction
Federated learning (FL) has attracted tremendous attention for executing data-intensive Internet-of-Things (IoT) applications (e.g., smart manufacturing), where data is distributedly collected at edge devices [1]. This distributed machine learning (ML) approach is an alternative to centralized ML methods that require transferring distributed data to cloud servers, which often cause a high communication overhead and concerns regarding the exposure of raw data to untrusted parties. FL performs ML training through the repetition of multiple global training rounds, with each round consisting of two key steps: (i) using their local data, clients/devices perform local model training, e.g., via several local iterations of stochastic gradient descent (SGD), and transfer their local models to a server, and (ii) the server aggregates all the received models (e.g., via averaging) to a global model and then broadcasts it back to the clients to initiate the next global training round.
@@ I-A Motivation
Despite its tremendous potential, FL is anticipated to be deployed on resource-constrained IoT wireless devices such as IoT sensors [1], which is challenging due to the increasing size of modern ML models that demand substantial computational power for local model training and incur high communication overheads during the transfer from devices to the central server. To this end, research – see [1] and the references therein – has focused on (i) tuning the number of local SGD iterations and frequency/period of client-to-server communications; (ii) model compression, gradient sparsification, or selective parameter sharing; (iii) efficient client selection/recruitment. Despite the significance of these methods, there are two key underlying assumptions made in all of them, limiting their applicability in real-world scenarios:
(A1) Snapshot-Based Analysis: These methods often assume that network and client conditions, such as client location and client-to-server channels, remain fixed during each global training round of FL, and thus they conduct snapshot-based analysis of the network. Afterwards, in each snapshot – often presumed to be the condition of the network at the beginning of each global training round – they make control decisions (e.g., spectrum allocation and client recruitment). However, in real-world scenarios, such analysis may quickly become sub-optimal, as the network condition may vary over time due to factors like client mobility and evolving client datasets.
(A2) Isolated Execution of a Single FL-Service: The existing methods primarily focus on executing a single FL-service. However, in large-scale IoT networks, multiple FL-services may simultaneously compete for wireless resources and client recruitment (e.g., Google may run an FL-service for next-word prediction, while Apple does so for face recognition). Despite its importance, only a few studies have explored this topic [2, 3], all of which consider snapshot-based analysis suffering from the above limitations.
@@
(2) Perspective-Driven Load Balancing: Load balancing refers to distributing the network traffic, preventing congestion across the network. The coexistence of multiple FL-services (relaxation of (A2)) calls for a unique load balancing strategy that considers a nuanced phenomenon which we call perspective-driven client distribution. This phenomenon captures the fact that the distribution of FL clients across different network regions varies depending on the perspective from which clients are evaluated/perceived. For example, regions with advanced connectivity infrastructure, such as urban areas, may have a higher density of clients suitable for latency-sensitive FL-services. Also, certain regions may have a higher density of low-cost clients, making them ideal for cost-sensitive FL-services, while other regions may have clients with greater computational capabilities, better suited for computation-intensive FL-services. Given the perspective-driven client distribution, in FL, load balancing must account for the overlap of the FL-services’ objectives.
@@ I-B Migration from Per-User to Service-Level Objectives
Simultaneous addressing of the above two challenges calls for client mobility management (i.e., handover design for load balancing) and dynamic resource allocation solutions, which are the key network control functionalities of radio access networks (RANs) [4, 5]. Additionally, orchestrating the coexistence of multiple FL-services over a dynamic IoT network, while addressing the above two challenges, demands a system management approach operating at different levels of agility. For example, while client recruitment for different FL-services can be performed at the start of each global round, mobility management requires a finer-granular time scale (real-time) to accommodate the fast-paced changes in network topology due to client repositioning. However, the existing IoT network control functionalities are often developed for native wireless services (e.g., high data-rate Internet provisioning) and primarily focus on per-user QoS requirements, including energy consumption and communication latency. Instead, in FL, addressing the aforementioned two challenges should consider collective service-level objectives/characteristics, such as the prediction accuracy of the FL models and client recruitment costs, alongside QoS metrics like overall energy consumption and model training latency across FL-services. Additionally, effective load balancing in FL must ensure an optimal allocation of clients to FL-services and wireless resources to the clients, while considering the innate differences among FL-services. For example, consider an FL-service for monitoring industrial equipments utilized during business hours coexisting with an FL-service focused on air quality monitoring. Equipment monitoring might require rapid ML model training to predict failures, while environmental sensing may tolerate longer ML training delays. In this case, the system should prioritize allocating network resources to industrial equipment monitoring during business hours and then shift resources towards environmental sensing.
@@
Figure 1: O-RAN architecture (all abbreviations follow standard 3GPP terminologies [6]. O-RAN components interact via standard open interfaces (E2, F1, open fronthaul, A1, and O1), facilitating interoperability between network elements from different manufacturers.
@@
These considerations necessitate a revamp of the existing IoT network control functionalities to facilitate the coexistence of multiple FL-services within a dynamic network – an approach absent in the literature. This naturally calls for a migration from traditional RANs (e.g., distributed RANs in 4G), operating with black-box network control functionalities applied to all network services [4]. To overcome this limitation and develop dedicated network control functionalities for FL-services, we leverage virtualization, intelligence, and programmability features of an emerging RAN technology used in 5G-and-beyond networks, called Open RAN.
@@
O-RAN (Fig. 1) transitions from rigid cellular to multi-vendor and data-driven networks by integrating the concepts of intelligence, virtualization (RAN slicing), and programmable white-box hardware (as opposed to the black-box hardware) [6]. O-RAN supports multiple services with diverse QoS requirements via RAN slicing, which partitions RAN resources into isolated RAN slices, each leased to a virtual network operator (VNO) [4]. Unlike traditional RANs, O-RAN fosters dedicated functionalities for each type of service, a feature referred to as programmability. For example, in contrast to the black-box MAC scheduler used in 4G that serves all services uniformly, O-RAN enables the implementation of a dedicated MAC scheduler for each service.
@@
In this work, we exploit O-RAN’s unique features, i.e., RAN slicing and programmability, and develop a novel intelligent FL architecture, called elastic FL (EFL). In EFL, we introduce dedicated network control functionalities to orchestrate multiple FL-services over a shared dynamic wireless network.
I-C Contributions
@@
  • In EFL, we use intelligence and programmability features of O-RAN to implement dedicated dynamic network control functionalities for FL-services (relaxation of (A1)) at three time scales: non-real-time (non-RT), near-RT, and RT.
@@ II EFL: Elastic FL
In the following, we systematically introduce the EFL framework, which will be implemented over O-RAN.
II-A Preliminaries on O-RAN
5G-and-beyond IoT networks are expected to support diverse services, including ultra-reliable low latency communications and massive machine-type communications [4]. They also are envisioned to provide services for various verticals – groups of companies requiring similar services (e.g., industrial factories) – operated by distinct virtual network operators. This versatile service provisioning is enabled by the use of intelligent and programmable O-RANs (Fig. 1). In O-RAN, 3GPP stack functionalities are disaggregated into (i) radio unit (O-RU), (ii) distributed unit (O-DU), and (iii) centralized unit (O-CU) [6]. O-RAN also introduces RAN intelligent controllers (RICs), i.e., non-real-time (non-RT) RIC and near-RT RICs, which utilize AI/ML algorithms to manage RAN operations at different time-scales [6, 7].
II-B Integration of EFL in O-RAN
Figure 2: Top Panel: The three main components of EFL (NSD, NFC, and FMS) are illustrated along with their key characteristics and distinguishing features. Bottom Panel: The process flow of EFL execution is depicted, showing the integration of NSD (red boxes), NFC (blue box), and FMS (green box). The sequence of operations is indicated by directional arrows, each annotated with a descriptive label written in blue text.
To integrate EFL in O-RAN, we first focus on intelligent network control in O-RAN. O-RAN supports intelligent control across three levels of agility: (i) Non-RT RIC operates on a timescale of seconds to minutes, establishing network management policies and hosting applications, known as rApps, which handle long-term functionalities like training AI models for network management. These trained AI models will be packaged into applications, referred to by xApps, and deployed on near-RT RIC. (ii) Near-RT RIC, operating on timescales of 10 milliseconds to seconds, leverages xApps to forecast traffic flow and network dynamics. Using these forecasts, near-RT RIC manages load balancing, network slicing, and mobility management. (iii) Finally, the MAC Scheduler, located at the O-DUs, operates in sub-10-millisecond intervals, focusing on real-time resource allocation to the clients.
EFL, depicted in Fig. 2, exploits O-RAN programmability and envisions dedicated functionalities at each of the aforementioned control layers to manage the concurrent execution of multiple FL-services (relaxation of (A2)) in a dynamic wireless environment (relaxation of (A1)). We refer to these functionalities as: (i) Non-RT System Descriptor, located at the non-RT RIC; (ii) Near-RT FL Controller, deployed at the near-RT RIC; (iii) FL MAC Scheduler, deployed at the O-DUs. We next design these functionalities to consider the collective service-level performance requirements, such as FL model accuracy across all FL-services.
II-C Non-RT System Descriptor
The coexistence of multiple FL-services introduces a new dimension of heterogeneity, which we call heterogeneity of interest. This heterogeneity reflects the diverse interests of FL-services in recruiting clients with different (i) datasets, (ii) data quality, (iii) channel conditions, and (iv) processing capabilities. These preferences are shaped by quality-of-service (QoS) requirements of each FL-service (e.g., target model accuracy and latency sensitivity) and its budget to recruit clients and acquire wireless resources.
@@ II-D Near-RT FL Controller
II-D1 FL slice creation
EFL enables the coexistence of multiple FL-services by creating a virtual RAN slice for each FL-service. In particular, using eApps, the near-RT FL controller first predicts the FL system states (e.g., client congestion at O-RUs). These predictions are derived from both current and historical data provided by the non-RT system descriptor. Based on these predictions, the near-RT FL controller then sends a slice creation request for each FL-service to the near-RT RIC responsible for RAN slicing. This request specifies (i) the required resources, such as bandwidth/spectrum, and (ii) the QoS requirements for each FL-service, including delay constraints and priority for accessing O-RAN resources. The slice is then provisioned with the wireless resources (e.g., spectrum) and sent to O-DUs where FL MAC schedulers perform real-time resource allocation for clients of FL-services.
@@ II-D1 FL slice creation
II-D2 Elastic slice scaling down/up
@@
II-D3 Client mobility management
@@
Figure 3: Real-world distribution of 150 FL clients over the Porto City Taxi Dataset: https://www.kaggle.com/datasets/crailtap/taxi-trajectory/data. Clients follow their actual mobility trajectories across the city and are served by $18$ O-RUs forming a realistic cellular network. These O-RUs are grouped under 44 O-DUs, such that each O-DU serves on average 4455 O-RUs. The colored dots in the figure represent the locations of individual clients, where different colors indicate clients managed by the same O-DU. The coverage areas of the O-RUs, each spanning 500 meters, are shown as colored hexagonal cells, where different colors indicate O-RUs managed by the same O-DU. In addition, the trajectories of the clients are shown by lines across the city. This setup creates a realistic dynamic wireless environment for evaluating slice management, mobility handling, and MAC scheduling within the proposed EFL framework.
Figure 4: The number of successful clients and its impact on the model prediction of FL-services under baseline methods and our proposed EFL method. Our proposed method outperforms the baseline methods in terms of the number of successful clients and the model prediction performance of FL-services. EFL leads to the success rate of above 0.9 for all the datasets, which is better than the best baseline (i.e., S-HMAC), which can reach the top success rate of 0.8. Additional insights can be found in the annotated text within the plots.
II-E FL MAC Scheduler
EFL introduces a new dimension to MAC schedulers, enabling client-level, real-time resource allocation for FL clients. This addresses resource over/under-provisioning at a shorter time-scale compared to near-RT FL controller, tailored to service-level performance metrics of FL-services, such as delay constraints for receiving local models from clients. Specifically, for FL-services, we equip EFL with a dedicated FL MAC scheduler located at O-DU to allocate wireless resources to FL clients.